# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 5. Main result: equivalence ${\mathfrak{F}}$
5.1. Functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$
Recall that in §2.3.3, we construct an algebra homomorphism
@@ 5.1. Functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$
As discussions in §4.2.2, the translation functor $T^{-\rho}_{0}:U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S}$ can be enhanced to a functor
@@
In fact, ${\mathbb{V}}_{r}$ is equivalent to the quotient functor $\bigoplus\limits_{\lambda\in\Lambda}\mathrm{Hom}(Q(w_{0}\bullet 0+l\lambda)_{S},-)$ (we will not use this below).
Recall that the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ is represented by the $G$-equivariant $\big{(}{\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}],{\mathrm{A}}\big{)}$-bimodule $\Gamma(\tilde{{\mathfrak{g}}},{\mathcal{E}})$. We can define a functor
${\mathbb{V}}_{c}:({\mathrm{A}}_{S}^{b}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$
as ${\mathbb{V}}^{b}_{c}$ in §4.2.2. It restricts to a functor
${\mathbb{V}}_{c}:{\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{B}\rightarrow\mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
@@
in $\mathrm{Coh}\big{(}({\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})_{\hat{0}}\big{)}$, where $\hat{0}$ represents the completion at $0\in{\mathfrak{t}}/W$. Tensoring with $B_{s}$ defines a functor on $\mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ and ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$.
It is automatic that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are compatible with $\Lambda$-translations. We now show that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are also compatible with reflection functors associated with $s\in{\mathrm{I}}$.
5.1.1. Compatibility with reflection functors
Let $\omega\in\Xi$. Consider the composition of functors
${\mathcal{O}}^{\omega}_{S}\xrightarrow{T_{\omega}^{0}}{\mathcal{O}}^{0}_{S}\rightarrow{\mathcal{O}}^{0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{T^{\omega}_{0}}{\mathcal{O}}^{\omega}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]},$
Lemma 5.1 ([37, Prop B.9]).
There is a natural isomorphism
$T_{0}^{\omega}\circ T_{\omega}^{0}\simeq-\otimes_{S^{W_{\omega}}}S$
@@ 5.1.1. Compatibility with reflection functors
$T^{-\rho}_{0}\circ\Theta^{r}_{s}$ $=T^{-\rho}_{s}\circ T^{s}_{0}\circ T^{0}_{s}\circ T_{0}^{s}$
$=T^{-\rho}_{s}\big{(}T_{0}^{s}(-)\otimes_{S^{s}}S\big{)}$
@@
where for the first and the last equality we use a natural isomorphism $T_{0}^{-\rho}=T_{s}^{-\rho}\circ T_{0}^{s}$. ∎
@@
Let $P$ be any (standard) parabolic subgroup of $G$, and let $W_{P}$ be the corresponding parabolic subgroup in $W$. Denote by $\pi^{P}:\tilde{{\mathfrak{g}}}_{P}\rightarrow{\mathfrak{g}}$ the natural projection. There is a ($G$-equivariant) natural isomorphism
$\pi_{P,*}\pi_{P}^{*}\simeq-\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W_{P}]}{\mathbb{C}}[{\mathfrak{t}}]\qquad\text{on}\quad D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{P}).$
@@
of functors from $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}})$ to $D^{b}\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$. Consider the special case when $P=P_{s}$, and then it implies our desired isomorphism by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$. The second assertion is easy to verify. ∎
@@ 5.2. Truncated categories
In this section, we define truncated subcategories in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ which will be shown to correspond to (certain) truncations in ${\mathcal{O}}^{0}_{S}$ defined in §2.3.1 under the equivalence (1.4), and we will study the projective objects in them.
We firstly define the truncation of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ by $\nu\in\Lambda$ as
${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}:=({\mathfrak{F}}^{b})^{-1}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0,\leq l\nu}_{S}).$
It is a Serre subcategory of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$, and each object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admits a finite decomposition with each direct factor contained in the truncated subcategory by some $\nu$. We define
${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu},\quad\nu\in\Lambda$
to be the full subcategory of modules in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ which are contained in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}$, forgetting the $B$-equivariant structure. Then ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ is a Serre subcategory in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$, and each object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admits a finite decomposition with each direct factor contained in the truncated subcategory by some $\nu$.
@@
Let $\leq^{\frac{\infty}{2}}$ be the semi-infinite order on $W_{\mathrm{af}}$, which is a partial order generated by
$xs\leq^{\frac{\infty}{2}}x\quad\text{if}\quad x\bullet_{l}0\geq xs\bullet_{l}0,\quad x\in W_{\mathrm{af}},\ s\in{\mathrm{I}}_{\mathrm{af}}.$

### v4

@@ 5. Main result: equivalence $\mathfrak{F}$}
5.1. Functors $\mathbb{V}$}
We construct the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$, and discuss their compatibility with reflection functors associated to $\mathrm{I}$.
5.1.1. Construction of $\mathbb{V}$}
Recall that in (2.8) we construct an algebra homomorphism
@@ 5.1.1. Construction of $\mathbb{V}$}
By similar discussions as in §4.4, the translation functor ${\mathsf{T}}^{-\rho}_{0}:U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S}$ can be enhanced to a functor
@@
Remark 5.1.
In fact, ${\mathbb{V}}_{r}$ is equivalent to the quotient functor $\bigoplus\limits_{\lambda\in\Lambda}\mathrm{Hom}(Q(w_{0}\bullet 0+l\lambda)_{S},-)$. We will not use this fact below.
Recall the functor ${\mathbb{V}}^{b}_{c}$ defined in (4.13). By $U{\mathfrak{n}}$-equivariance it upgrades to a functor
${\mathbb{V}}_{c}:\ ({\mathsf{A}}_{S}^{b}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T},$
which restricts to a functor
${\mathbb{V}}_{c}:\ {\mathsf{A}}_{S}^{b}\text{-}\mathrm{Mod}^{B}\rightarrow\mathrm{QCoh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
It is automatic that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are compatible with $\Lambda$-translations.
5.1.2. Compatibility with reflection functors
We show that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are compatible with reflection functors associated with $s\in{\mathrm{I}}$. The results of this subsection will not be used in the sequel.
Let $\omega\in\Xi$. Consider the composition of functors
${\mathcal{O}}^{\omega}_{S}\xrightarrow{{\mathsf{T}}_{\omega}^{0}}{\mathcal{O}}^{0}_{S}\rightarrow{\mathcal{O}}^{0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{{\mathsf{T}}^{\omega}_{0}}{\mathcal{O}}^{\omega}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]},$
Lemma 5.2 ([35, Prop B.9]).
There is a natural isomorphism
${\mathsf{T}}_{0}^{\omega}\circ{\mathsf{T}}_{\omega}^{0}\simeq-\otimes_{S^{W_{\omega}}}S$
@@ 5.1.2. Compatibility with reflection functors
[…]
@@
${\mathsf{T}}^{-\rho}_{0}\circ\Theta^{r}_{s}$ $={\mathsf{T}}^{-\rho}_{s}\circ{\mathsf{T}}^{s}_{0}\circ{\mathsf{T}}^{0}_{s}\circ{\mathsf{T}}_{0}^{s}$
$={\mathsf{T}}^{-\rho}_{s}\big({\mathsf{T}}_{0}^{s}(-)\otimes_{S^{s}}S\big)$
@@
where for the first and the last equality we use a natural isomorphism ${\mathsf{T}}_{0}^{-\rho}={\mathsf{T}}_{s}^{-\rho}\circ{\mathsf{T}}_{0}^{s}$. ∎
@@
Let $P$ be any (standard) parabolic subgroup of $G$, and let $W_{P}$ be the corresponding parabolic subgroup in $W$. Denote by $\pi^{P}:\widetilde{{\mathfrak{g}}}_{P}\rightarrow{\mathfrak{g}}$ the natural projection. There is a ($G$-equivariant) natural isomorphism
$\pi_{P,*}\pi_{P}^{*}\simeq-\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W_{P}]}{\mathbb{C}}[{\mathfrak{t}}]\qquad\text{on}\quad D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{P}).$
@@
[…]
@@ 5.2. Truncated categories
In this subsection, we review the truncated subcategories of ${\mathcal{O}}^{0}_{S}$ studied in [34, §4.2], which is a refinement of the truncations discussed in §2.3.1. Next, we define the corresponding truncated subcategories in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$. The advantage of the new truncations is their compatibility with the functors $\mathbb{V}$}, see Lemmas 5.10 and 5.19.
5.2.1. Truncation in ${\mathcal{O}}^{0}_{S}$
Definition 5.6.
@@ 5.2.1. Truncation in ${\mathcal{O}}^{0}_{S}$
The semi-infinite order $\leq^{\frac{\infty}{2}}$ on $W_{\mathrm{ex}}$ is the partial order generated by
$x\leq^{\frac{\infty}{2}}tx\quad\text{if}\quad x\bullet_{l}0\leq tx\bullet_{l}0,$
for $x\in W_{\mathrm{ex}}$ and reflection $t\in\bigcup_{x\in W_{\mathrm{af}}}x{\mathrm{I}}_{\mathrm{af}}x^{-1}$ in $W_{\mathrm{af}}$.
Remark 5.7.
  1. (1)